OpenStudy (anonymous):

Function Factors f(x) = 3x3 + x2 − 38x + 24 (x + 4), (x − 3) How do I find the remaining factors of f? this is pre calculus btw

OpenStudy (anonymous):

divide is one way, or reason it out is the other

OpenStudy (anonymous):

\[(x+4)(x-3)=x^2+x-12\] you want \[ 3x^3 + x2 − 38x + 24 \] so you need \(ax+b)\) where \(ax\times x^2=3x^2\) and so \(a=3\)

OpenStudy (anonymous):

also you need \(b\times(-12)=24\) so \(b=-2\)